A person climbs to the top of a mountain at the rate of 1.5 km/h and down the same way at a rate of 4.5 km/h. If he is gone for 6 hours, how far did he travel from camp to the top of the mountain?
Answer by Alan3354(69443)

A person climbs to the top of a mountain at the rate of 1.5 km/h and down the same way at a rate of 4.5 km/h. If he is gone for 6 hours, how far did he travel from camp to the top of the mountain?
----------
Avg speed of the round trip = 2*1.5*4.5/(1.5+4.5) = 13.5/6 = 2.5 km/hr
---
2.5*6 = 15 km round trip
--> 7.5 km each way